1 mole of ferric oxalte oxidised completely by x mol of MnO4- in acidic medium, then x is
 

Grade:12th pass

1 Answers

Vikas TU
7 years ago
From simple mole concept -2 moles x valence factor = contant we can calculate.
.i.e. 
ferric oxalate and MnO4- would have equal equivalents.
Therefore,
1 * 2 = x * 5
x = 2/5 = > 0.4 mol